# May I Ask a Follow-up Question? Understanding the Benefits of Conversations in Neural Network Explainability

### v1

@@ 2.3. Human-AI Collaboration
Human-AI collaboration is an emerging research area, which explores how humans and AI systems can work together to achieve shared goals (Arous et al., 2020; Lai et al., 2022; Kim et al., 2023). Prior studies within this domain have investigated collaborations between humans and various AI systems, from robots (Gero et al., 2020; Häuslschmid et al., 2017; Kumar et al., 2021; Ogenyi et al., 2021) to virtual agents (Bickmore et al., 2009; Ashktorab et al., 2020; Cai et al., 2019; D’Avella et al., 2022; Numata et al., 2020). The tasks involved span a broad scope, including text (Bansal et al., 2021) and image (Kim et al., 2023) classifications, medical diagnosis (Cai et al., 2019), deception detection (Lai and Tan, 2019) and cooperative games (Gero et al., 2020; Ashktorab et al., 2020; Feng and Boyd-Graber, 2019). An area of particular interest within these collaborations is the role of explanations in influencing human-AI decision-making (Nguyen et al., 2021, 2022; Lai and Tan, 2019; Bansal et al., 2021).
Our study aligns with existing work on human-AI collaboration (Nguyen et al., 2021, 2022; Lai and Tan, 2019; Bansal et al., 2021; Feng and Boyd-Graber, 2019). In our work, participants need to collaborate with explanations to choose the most accurate neural networks among others. Instead of exploring the role of explanations in collaboration, we mainly examine the potential of conversations in aiding users to effectively collaborate with explanations.
@@
3. Methodology
Our study aims to investigate the impact of conversations on the explainability of AI models by observing participants’ comprehension, acceptance, trust of the static explanations, and ability to use the explanations to select the most accurate neural networks before and after the conversation. We devised both objective and subjective evaluations. For the objective evaluation, participants are presented with three neural networks and need to identify the most accurate one based on the provided explanations. All three networks make the same decisions on several images; the only differences between the candidates are the explanations of their predictions. Consequently, to make the correct choice, participants must understand the explanations. Additionally, we gauge participants’ perceptions of static explanations, including understanding, acceptance, and trust. This section delves into a comprehensive overview of our research methodology. Our study has received approval from the Institutional Review Board at the University (#IRB-2023-254).
3.1. Task, AI models and Explanation Method Selection
In our study, we focus on the image classification task — a cornerstone in the field of computer vision (CV) that has been the subject of various human-AI collaborative studies (Taesiri et al., 2022; Jeyakumar et al., 2020). Image classification, by virtue of its minimal domain-specific expertise requirement, lends itself well to crowdsourcing where participants are from diverse domains. We specifically choose the ImageNet dataset (Deng et al., 2009) for this purpose, which boasts 14,197,122 annotations aligned with the WordNet hierarchy and stands as a significant benchmark in image classification and object detection. In our experiments, we select and train three classification models, each exhibiting varied top-1 accuracies: the Swin Transformer (84.1%) (Liu et al., 2021b), VGG-16 (71.6%) (Simonyan and Zisserman, 2015), and AlexNet (56.5%) (Krizhevsky et al., 2012).  In the objective evaluation, participants need to choose, from these three classification models, the best one based on provided explanations (detailed in Section 3.2.1). Alongside the conversation panel in the chat interface (Fig. 3), we display the predictions and the explanations of only the Swin Transformer as a reminder of the task. But participants can still ask questions about any other classification models.
We focus our research on feature attribution explanation methods, which generate explanations by investigating the importance of different input features to final predictions. The reason is that feature attribution explanations are more intuitive to understand for laypeople as they directly show the relationship between input features and model predictions (Kim et al., 2023). Specifically, we select two explanation techniques from two main categories of feature attribution explanation methods: LIME (Ribeiro et al., 2016) (surrogate methods) and Grad-CAM (Selvaraju et al., 2017) (gradient-based methods). LIME works by training an inherently interpretable surrogate model, e.g., a linear model, around a specific prediction. It uses the simple surrogate model to explain the predictions from complex classification models. It spotlights the most crucial features of a model input. Grad-CAM manifests its explanations via a heatmap, color-coding regions of an input image based on their significance to the model’s prediction. It works by calculating the gradients of the output class with respect to the feature maps in the final convolutional layer of a deep neural network and using these gradients to weight the feature maps. The resulting weighted feature maps are then averaged to generate a heatmap that highlights the most activated regions of the image. Figure 2 displays example explanations generated by these two explanation methods. Participants are randomly assigned with an explanation method to do the study once they log into our system.
3.2. Study Design
We devise both subjective and objective evaluations to assess the effect of conversations on participants’ understanding, acceptance, and trust of static explanations. We conduct these evaluations both before and after conversations. By comparing the results, we estimate the impact on conversations for static explanations.
@@ 3.2. Study Design
Comprehension
How much do you think you understand the explanations provided for predictions of deep learning models?
Perceived Usefulness
Using explanations would improve my understanding of deep learning models’ predictions.
@@
Perceived Ease-of-Use
I become confused when I use the explanation information.
@@
Behavioral Intention
I would prefer getting explanation information as long as it is available when getting predictions from deep learning models.
@@
Trust
How would you rate the competence of the explanation method? - i.e. to what extent does the explanation method perform its function properly?
@@
Table 1. Detailed questions in the subjective evaluation, which measures participants’ feelings about static explanation from five aspects, i.e., Comprehension, Perceived Usefulness, Perceived Ease-of-Use, Behavioral Intention, and Trust. The user will respond to each question using a 7-point Likert scale.
3.2.1. Objective Evaluation – Decision-Making of Classification Models
@@
The evaluation aims to objectively evaluate participants’ understanding of static explanations. We ask participants to choose, from three classification models, the most accurate on unobserved test data. The three classification models yield identical decisions on 5 images and come with static explanations from the same method. The only differences between the three networks lie in their explanations. Hence, to make the correct selection, the participants must understand the explanations.
We choose three classification models with different performances for this evaluation: Swin Transformer with 84.1% accuracy, VGG-16 with 71.6% accuracy, and AlexNet with 56.5% accuracy. Figure 2 presents a sample image for this evaluation. The full set of images can be accessed in Appendix A. For evaluation integrity, the order of different classification models is random for each participant and the examples used in the evaluation differ from those displayed during conversations.
To ensure that users can make the correct choice based on the explanations, it is crucial that the explanations for model predictions provide clear indicators of each model’s performance. While we recognize that high-performing classification models generally produce more reasonable explanations than low-performing ones, this is not always true for two main reasons. Firstly, the accuracy of a model does not guarantee the plausibility of its explanations. A lower-performing model might align more with human intuition and thus generate seemingly reasonable explanations. Secondly, as we choose examples where all three models make the same predictions, their explanations might be all the same and reasonable. Therefore, we have selected explanation examples where better classification models indeed provide more reasonable explanations. This approach allows users to evaluate and select the best classification models based solely on the explanations, thus making their choice a true reflection of their objective understanding.
An explanation is deemed more reasonable when it focuses on more features that are unique to the predicted class. For example, an explanation that highlights both the head and tail of a fish is considered better than an explanation focusing on the fish head only. This is because a classification model that relies on multiple features for predictions is likely to still make the correct decision if a subset of features are missing or obscured in the image. In the example presented in Figure 2, Model is better than Model A or Model C as Model B utilizes both the head and the body for classification, and does not focus on the background like Model A.
@@ 3.2.2. Subjective Evaluation
We also measure participants’ subjective perceptions of static explanations, including their comprehension, acceptance, and trust. Each question is chosen based on an in-depth review of existing literature and has been validated in prior research. Subjective Evaluation contains a total of 13 questions, all of which utilize a 7-point Likert scale for responses. Table 1 lists all questions used in the subjective evaluation.
@@
  • Trust (Muir and Moray, 1996): Participants’ confidence in the reliability of the explanation methods to function as intended. Trust has been recognized as an important factor in human-AI collaboration as it mediates the human’s reliance on AI models, thus directly affecting the effectiveness of the human-AI team (Sebo et al., 2020; Seaborn et al., 2021; Doshi-Velez and Kim, 2017). Recent studies show that increased transparency in ML systems doesn’t significantly impact people’s trust (Cheng et al., 2019; Schaffer et al., 2019; Kunkel et al., 2019; Poursabzi-Sangdeh et al., 2021), possibly due to complex trust mechanisms (McKnight et al., 1998; Zhang et al., 2020) and potentially overloaded information (Schaffer et al., 2019; Poursabzi-Sangdeh et al., 2021; Feldhus et al., 2022). In this study, our intention is to discern if clarifying participants’ confusion via conversations about explanations can influence their trust in explanations.
Discipline Number of Participants
Business
8
Psychology
7
Accountancy
6
Biology
6
Economics
5
Chemistry, Chemical Engineering and Biotechnology
4
Sociology
4
Communication Studies
3
Linguistics and Multilingual Studies
3
Civil and Environmental Engineering
2
Computer Science
2
Electrical and Electronics Engineering
2
Sport Science & Management
2
Design (Architecture)
1
Foods science
1
Mathematics
1
Medicine
1
Maritime Studies
1
Physics
1
Table 2. Disciplines of our participants and the number of participants in each discipline. There are 60 participants from 19 different disciplines in total.
3.3. Study Procedure
@@
Before participation, individuals are required to sign an informed consent form that outlines the study’s objectives and procedures. The form also clarifies compensation details and assures both the anonymity and confidentiality of data collected during the study. Upon signing the consent, participants receive an email providing a website link and login credentials (username and password). They are instructed to use these credentials to access the study platform. Details about the study platform are in section 3.6.
@@
After logging in, a pop-up prompt provides an overview of the tasks ahead. Participants then complete pre-conversation objective and subjective evaluations of static explanations. The objective evaluation measures participants’ understanding of static explanations by letting them choose, from three classification models, the most accurate on unobserved test data. The subjective evaluation, with 13 self-reporting questions, probes their perceived comprehension, acceptance, and trust towards the static explanations. Following these evaluations, participants engage in a discussion about static explanations with an XAI expert to deepen their understanding. During the conversation, the example images displayed are different from those used in the evaluations; however, the explanation methods remain the same. Participants are motivated to understand the explanations as they need to select the best-performing classification model using explanations only when doing objective evaluation. Our XAI experts faithfully answer the user’s questions based on their knowledge, trying to help the user gradually understand the explanation. After the conversation, participants complete the same set of evaluations as before. All evaluation outcomes and conversation records are documented. Upon study completion, each participant receives a \$10 reward.
3.4. Wizard-of-Oz Experimental Setting
We conduct conversational explanations with participants in the Wizard-of-Oz (WoZ) setting (Kelley, 1984). Participants interact with a dialogue system that they believe to be autonomous but is actually operated by a human being, the Wizard. The reason for using WoZ is that we can use the collected conversations for future dialogue system development. Several works have shown that the WoZ approach can be applied to collect high-quality typed conversations where a machine learning-based system can learn from(Wen et al., 2017; Budzianowski et al., 2018; El Asri et al., 2017).
During conversations, participants are encouraged to pose questions or offer comments on the provided static explanations. To maintain consistency in the wizard responses, we develop a question-answer repository. We populate an initial repository based on the understanding of the explanation methods (Selvaraju et al., 2017; Ribeiro et al., 2016) and a pilot study with 6 participants. During the actual experiment, when a participant presents a new question, the corresponding answer is added to the question-answer repository. If the same question occurs again, we refer to the previously stored answer. Section 4.5 presents further discussion of the questions from participants during our study.
Figure 3. The Conversation page where users can discuss static explanations with an expert.
3.5. Participant Recruitment and Selection
We recruit our participants in two ways: by posting advertisements on an online forum and by emailing students across various departments and schools. To join our study, participants have to be 21 years old or older, fluent in English, and not involved in research about XAI. To avoid bias in the study and mirror real-world utilization of XAI, we selectively recruit participants to promote the diversity in themajors / discipline of specialization of the participants. Our study is conducted remotely. To ensure data quality, participants are excluded from data analysis if they are away from the keyboard for more than 10 minutes during the study. This results in a total of 60 participants from 19 majors. Table 2 shows statistics of the majors that the participants are enrolled in. Previous research has shown that repeated interactions with AI-based decision-making systems can affect decision subjects’ fairness perceptions (Gemalmaz and Yin, 2022). To mitigate potential bias, each participant can partake in our study only once.
3.6. User Interface design
To conduct the study, we design and build a web-based platform where participants can remotely finish the whole procedure of the experiment. The web-based platform will ensure that all communications between users and XAI experts are text-based and recorded. Our platform will contain three different web pages, i.e., the Login page, the Conversation page, and the Evaluations page. The Login page allows users to log into our platform by using the credentials (username and password) provided to them. After users log onto the platform, they are directed to the Evaluations page. The Evaluations page helps participants complete objective and subjective evaluations (more details in section  3.2.1 and 3.2.2): choosing the most accurate model based on explanations and reporting their perceptions of static explanations. Pre-conversation evaluations are conducted before starting conversations and post-conversation evaluations are conducted after they discuss static explanations with XAI experts. The Conversation page is where participants discuss static explanations with XAI experts. There are two sections on the page. On the left, the user sees a task description, a description of the prediction model, a description of an explanation technique, a model input, a model output, an explanation from the explanation model, and a brief description of the explanation. Within the chatbox on the right, the user can converse with an XAI expert to clarify the explanation. Through a conversation, a user can ask any questions or provide any comments related to the explanation on the left side and our XAI experts will faithfully answer his/her questions based on their knowledge, trying to help the user gradually understand the explanation. Figure 3 shows an example screenshot of the Conversation page.
@@ 3.6. User Interface design
Explanation Methods Evaluation Timing Objective Understanding (Decision-Making Accuracy) Subjective Understanding Perceived Usefulness Perceived Ease of Use Behavioral Intention Trust
LIME before 0.38 $\pm$ 0.20 4.03 $\pm$ 1.35 5.09 $\pm$ 1.07 4.48 $\pm$ 0.94 5.25 $\pm$ 0.95 4.15 $\pm$ 0.88
after $*{}^{*}$$\pm$ 0.16 $*{}^{*}$$\pm$ 0.88 $*{}^{*}$$\pm$ 0.66 $*{}^{*}$$\pm$ 0.84 $*{}^{*}$$\pm$ 0.81 $*{}^{*}$$\pm$ 0.73
Grad-CAM before 0.82 $\pm$ 0.21 4.17 $\pm$ 0.91 5.49 $\pm$ 0.97 4.71 $\pm$ 0.95 5.52 $\pm$ 0.65 4.40 $\pm$ 1.00
after $*{}^{*}$$\pm$ 0.11 $*{}^{*}$$\pm$ 0.97 $*{}^{*}$$\pm$ 0.60 $*{}^{*}$$\pm$ 0.82 $*{}^{*}$$\pm$ 0.79 $*{}^{*}$$\pm$ 0.80
Table 3. Results of objective evaluation (decision-making accuracy), self-reporting comprehension score, perceived usefulness, perceived ease of use, behavioral intention, and trust of different static explanations before and after conversations. Each score is presented as mean $\pm$ standard deviation. $*{}^{*}$ indicates that post-conversation scores are statistically higher than pre-conversation scores with $p<0.001$ using the Student’s t-test.
(a) Results for LIME
(b) Results for Grad-CAM
Figure 4. Results of objective evaluation (decision-making accuracy), participants’ subjective understanding, perceived usefulness, ease of use, behavioral intention, and trust in static explanations, before and after conversational explanations. Decision-making accuracy is ranged from 0 to 1 and the rest scores are from 1 to 7. Figure (a) displays the scores for LIME and Figure (b) displays the average scores for Grad-CAM.

### v2

@@ 2.3. Human-AI Collaboration
Human-AI collaboration is an emerging research area, which explores how humans and AI systems can work together to achieve shared goals (Herse et al., 2023; Xu et al., 2023; Kim et al., 2023). Prior studies within this domain have investigated collaborations between humans and various AI systems, from robots (Gero et al., 2020; Häuslschmid et al., 2017; Carissoli et al., 2023; Liu et al., 2022; Bhat et al., 2024) to virtual agents (Ashktorab et al., 2020; Cai et al., 2019; D’Avella et al., 2022; Numata et al., 2020). The tasks involved span a broad scope, including text (Bansal et al., 2021) and image (Kim et al., 2023) classifications, medical diagnosis (Cai et al., 2019), deception detection (Lai and Tan, 2019) and cooperative games (Gero et al., 2020; Ashktorab et al., 2020; Feng and Boyd-Graber, 2019). An area of particular interest within these collaborations is the role of explanations in influencing human-AI decision-making (Nguyen et al., 2021, 2022; Lai and Tan, 2019; Bansal et al., 2021).
Our study aligns with existing work on human-AI collaboration (Nguyen et al., 2021, 2022; Lai and Tan, 2019; Bansal et al., 2021; Feng and Boyd-Graber, 2019). In our work, participants need to collaborate with explanations to choose the most accurate neural networks among others. Instead of exploring the role of explanations in collaboration, we mainly examine the potential of conversations in aiding users to effectively use explainability techniques and understand their outputs.
3. Method
Our study aims to investigate the impact of conversations on the explainability of AI models by observing participants’ comprehension, acceptance, trust of the static explanations, and ability to use the explanations to select the most accurate neural networks before and after the conversation. Our study has received approval from the Institutional Review Board at Nanyang Technological University (#IRB-2023-254).
Table 1. Academic disciplines of our participants and the number of participants in each group. There are 120 participants from 4 different discipline groups.
Academic Discipline Number of Participants
Business 23
Engineering 16
Humanities 55
Science 26
3.1. Participants
A total of 120 participants joined our study. All were 21 years old or older, fluent in English, and had not been involved in research about XAI previously. We recruited our participants in two ways: by posting advertisements on an online forum and by emailing students and staff across various departments and schools. They are from a wide range of disciplines to promote diversity. For ease of reporting, we categorize their disciplines into four groups:
  • Business, including Business and Accountancy.
  • Engineering, including Civil and Environmental Engineering, Computer Science, Electrical and Electronics Engineering, Maritime Studies, and Food Science.
  • Humanities, including Psychology, Economics, Communication Studies, Linguistics and Multilingual Studies, and Sociology.
  • Science, including Biology, Chemistry, Chemical Engineering and Biotechnology, Sport Science & Management, Mathematics, Medicine, and Physics.
Table 1 shows statistics of the academic disciplines that the participants enrolled in.
3.2. Experimental Task
@@
In our study, we focus on the image classification task on the ImageNet dataset (Deng et al., 2009). Image classification task is a cornerstone in the field of computer vision (CV) that has been the subject of various human-AI collaborative studies (Taesiri et al., 2022; Jeyakumar et al., 2020). We train three classification models with different top-1 classification accuracies: Swin Transformer (Liu et al., 2021b) (84.1%), VGG-16 (Simonyan and Zisserman, 2015) (71.6%), and AlexNet (Krizhevsky et al., 2012) (56.5%). To generate explanations for model predictions, we select two explanation techniques from two main categories of feature attribution explanation methods: LIME (Ribeiro et al., 2016) (a surrogate method) and Grad-CAM (Selvaraju et al., 2017) (a gradient-based method). We focus on feature attribution explanations as we believe the relationship between input features and model predictions is more intuitive to understand than example-based methods for laypeople (Kim et al., 2023). Figure 1 displays example explanations generated by these two explanation methods.
@@
To conduct the study, we design and build a web-based platform where participants can remotely finish the whole procedure of the experiment. After users log into the platform, we first evaluate their objective and subjective understanding of static explanations. The objective explanations require participants to choose, from three classification models, the most accurate on unobserved test data. The three classification models yield identical decisions on 5 images. The only differences between the three networks lie in their explanations. Hence, to select the best model, the participants must rely on the explanations. Figure 2 presents an example question, including the original image, the model outputs, and the explanations. The full set of questions used in the study can be found in Appendix A.
Table 2. Detailed questions in the subjective evaluation. The user will respond to each question using a 7-point Likert scale.
@@
Comprehension How much do you think you understand the explanations provided for predictions of deep learning models?
Perceived Usefulness Using explanations would improve my understanding of deep learning models’ predictions.
@@
Perceived Ease-of-Use I become confused when I use the explanation information.
@@
Behavioral Intention I would prefer getting explanation information as long as it is available when getting predictions from deep learning models.
@@
Trust How would you rate the competence of the explanation method? - i.e. to what extent does the explanation method perform its function properly?
@@
The subjective evaluation measures participants’ self-reported perception of the static explanations, including their comprehension (Hoffman et al., 2018; Cheng et al., 2019), acceptance (Davis, 1989; Davis et al., 1989; Diop et al., 2019; Flathmann et al., 2023), and trust (Davis, 1989; Davis et al., 1989; Diop et al., 2019; Yang et al., 2017b; Guo et al., 2023). Based on an in-depth review of existing literature, we chose the questions from those that have been validated in prior research. The subjective evaluation contains a total of 13 questions, each utilizing a 7-point Likert scale for responses. Table 2 lists all the questions we used. Labels of the 7-point Likert scale are listed in Appendix A.
Figure 3. The web page where users can discuss static explanations with an expert.
After these two evaluations, participants are divided into two groups, i.e., the control group and the experimental group. Participants in the control group read static explanations for 15 minutes. Participants in the experimental group conduct conversational explanations with participants in the Wizard-of-Oz (WoZ) setting (Kelley, 1984). They interact with a dialogue system that they believe to be autonomous but is actually operated by a human expert on machine learning.
To support the WoZ experiment, we built a conversation page with a two-section structure, as depicted in Figure 3. On the left, the page shows a task description, a textual description of the prediction model, a textual description of the explanation technique, an example input image, the model prediction on the input image, a static explanation for the prediction, and a textual description of the explanation. On the right, the interface enables users to converse with XAI experts, seeking clarifications and posing questions about the explanation. For the users in the control group, we replace the textual chat user interface with a 15-minute timer. Once the timer reaches zero, users are allowed to proceed to the post-evaluations. Users from both groups receive the same post-evaluations, which are identical to the pre-evaluations. We discuss the evaluations below.
3.3. Experimental Design
There are two independent variables and two categories of dependent variables. The independent variable in the experiments is the explanation method: LIME or Grad-CAM and the method of understanding static explanations: conversation with human experts or reading static explanations. As we devise both subjective and objective evaluations before and after conversations or readings, two categories of dependent variables were collected in the experiment: the model selection accuracy and the self-reported perception scores.
3.3.1. Objective Evaluation – Selection of Classification Models
@@
The evaluation aims to objectively evaluate participants’ understanding of the static explanations. Participants are presented with 5 input images, on which the three neural networks make the same decisions. The only differences between the three networks lie in their explanations. Participants need to choose the one that would be the most accurate on unobserved test data. Hence, to make the correct selection, the participants must understand the explanations. We use the accuracy of selecting the correct model to measure participants’ objective understanding of static explanations.
We recognize that existing explanation techniques are not always faithful to the underlying model (Adebayo et al., 2018; Kindermans et al., 2019; Jacovi and Goldberg, 2020) and do not always provide actionable information for model selection. As our goal is to test if the users can understand the static explanations when they do provide actionable information, rather than evaluating the static explanations themselves, we selected input images where better classification models indeed have more reasonable and intuitive explanations. This approach allows users to easily pick the best classification models if they understand the static explanations well. We deem an explanation more reasonable when it focuses more on discriminative features that are unique to the predicted class and less on spurious features that are irrelevant to the class. In addition, good models should have explanations that rely on multiple types of discriminative features. This is because a model relying on multiple features is robust and makes the correct decision even if some discriminative features are missing or occluded. In the example in Figure 2, Model B is better than Model A or Model C as Model B utilizes both the head and the body of the cat for classification. In addition, unlike Model A, Model B does not focus on the background, which is irrelevant to the predicted class, Siamese Cat.
@@ 3.3.2. Subjective Evaluation
We also measure participants’ subjective perception of static explanations, including their comprehension, acceptance, and trust. The subjective evaluation contains a total of 13 questions listed in table 2. All questions utilize a 7-point Likert scale for responses.
@@
  • Trust (Muir and Moray, 1996; Bach et al., 2022): Participants’ confidence in the explanation methods keeping functioning as intended. Trust has been recognized as an important factor in human-AI collaboration as it mediates the human’s reliance on AI models, thus directly affecting the effectiveness of the human-AI team (Sebo et al., 2020; Seaborn et al., 2021; Doshi-Velez and Kim, 2017; Vorm and Combs, 2022; Silva et al., 2023).
    The literature demonstrated that static explanations have inconsistent effects on users’ trust in AI systems. On one hand, several studies have demonstrated that detailed explanations (Glass et al., 2008; Ha and Kim, 2023; Silva et al., 2023), contrastive explanations (Larasati et al., 2020), and example-based explanations (Yang et al., 2020) can enhance user trust in systems. On the other hand, studies showed that static explanations do not have strong effects on user trust in AI systems (Cheng et al., 2019; Kunkel et al., 2019; Wang and Yin, 2021; Zhang et al., 2020).
    One main reason for these inconsistent reports is that trust is mediated by the users’ understanding of the static explanations (Kunkel et al., 2019; Wang and Yin, 2021; Zhang et al., 2020), and such understanding is often absent. According to theories of trust (McKnight et al., 1998; Lim et al., 2009; Hoffman et al., 2018), the ability to build a mental model of AI systems is the key for user trust in AI. Unsurprisingly, studies on the effects of static explanations for laypersons show that users with limited knowledge of machine learning struggle to understand static explanations and the decision-making processes they are supposed to explain. Consequently, these users do not exhibit increased trust in AI systems after receiving static explanations (Zhang et al., 2020; Wang and Yin, 2021).
    With this paper, we quantitatively investigate whether customized conversations about static model explanations can enhance user understanding and improve trust. The conversational approach toward explanations has been advocated by previous studies (Glass et al., 2008; Pieters, 2011; Schaffer et al., 2019; Feldhus et al., 2022; Lakkaraju et al., 2022) but never experimentally verified. For example, through interviews with decision-makers, Lakkaraju et al. (2022) found that decision-makers strongly prefer conversational explanations that allow them to ask follow-up questions.
3.4. Detailed Study Procedure
@@
Before participation, individuals are required to sign an informed consent form that outlines the objectives and procedures of the study. The form also clarifies compensation details and assures both the anonymity and confidentiality of data collected during the study. Upon signing the consent, participants receive an email that guides them to access the study platform.
@@
After logging in, a pop-up prompt provides an overview of the tasks ahead. Participants then complete pre-experiment objective and subjective evaluations of the static explanations. The objective evaluation measures participants’ understanding of static explanations by letting them choose, from three classification models, the most accurate on unobserved test data. There are 5 explanation examples in the objective evaluation. The subjective evaluation, with 13 self-reporting questions, probes the perceived comprehension, acceptance, and trust towards the static explanations. Following these evaluations, participants in the experimental group engage in a WoZ discussion about static explanations. During the conversation, one example image is displayed on the screen. The example image is different from those used in the evaluations; however, the explanation methods remain the same. Participants are motivated to understand the explanations as they need to select the best-performing classification model using explanations only when doing objective evaluation. Our XAI experts faithfully answer the user’s questions based on their knowledge, trying to help the user gradually understand the explanation. For participants in the control group, they read the static explanation for 15 minutes which is the average conversation time of the experimental group. After the conversation or 15-minute reading, participants complete the same set of evaluations as before. All evaluation outcomes and conversation records are documented. Upon study completion, each participant receives a \$10 reward.
Table 3. Results of the experimental group before and after conversations, and the control group before and after 15-minute reading. Each score is presented as mean $\pm$ standard deviation and the change $\delta$ before and after. $p<0.001$
@@
Explanation Methods Group Evaluation Timing Objective Understanding (Decision-Making Accuracy) Subjective Understanding Perceived Usefulness Perceived Ease of Use Behavioral Intention Trust
LIME experimental before 0.38 $\pm$ 0.20 4.03 $\pm$ 1.35 5.09 $\pm$ 1.07 4.48 $\pm$ 0.94 5.25 $\pm$ 0.95 4.15 $\pm$ 0.88
after 0.53 $\pm$ 0.16 5.30 $\pm$ 0.88 5.92 $\pm$ 0.66 5.28 $\pm$ 0.84 5.83 $\pm$ 0.81 4.92 $\pm$ 0.73
control before 0.37 $\pm$ 0.17 4.57 $\pm$ 1.43 5.67 $\pm$ 0.95 4.87 $\pm$ 1.26 5.73 $\pm$ 0.69 4.37 $\pm$ 0.90
after 0.40 $\pm$ 0.20 4.60 $\pm$ 1.16 5.33 $\pm$ 0.96 4.48 $\pm$ 1.26 5.27 $\pm$ 1.08 4.36 $\pm$ 1.05
Grad-CAM experimental before 0.82 $\pm$ 0.21 4.17 $\pm$ 0.91 5.49 $\pm$ 0.97 4.71 $\pm$ 0.95 5.52 $\pm$ 0.65 4.40 $\pm$ 1.00
after 0.92 $\pm$ 0.11 5.43 $\pm$ 0.97 6.12 $\pm$ 0.60 5.58 $\pm$ 0.82 6.08 $\pm$ 0.79 5.19 $\pm$ 0.80
control before 0.81 $\pm$ 0.20 4.07 $\pm$ 1.34 5.58 $\pm$ 0.59 4.36 $\pm$ 1.15 5.45 $\pm$ 0.71 4.22 $\pm$ 0.96
after 0.79 $\pm$ 0.19 4.40 $\pm$ 1.28 5.46 $\pm$ 0.69 4.70 $\pm$ 1.21 5.33 $\pm$ 0.83 4.42 $\pm$ 0.87